Shack Diesel Company has stock trading today at $40.59 per share. The company just paid a $2.46 per share dividend. Kenny Smith, an analyst, believes investors should seek a 10.00% return to hold the stock. If we value Shack Diesel using the constant growth model, what dividend growth rate would justify these assumptions?
Answer #1
P0 = D0*(1+g) / ( r - g )
Here,
P0 = Current value of stock = 40.59
D0 = Current dividend = 2.46
g = Growth rate
r = Required cost of capital = 10%
Now let us put these figures in the above equation to find out required return.
40.59 = 2.46 * ( 1 + g ) / ( 10% - g )
40.59 / 2.46 = ( 1 + g ) / ( 10% - g )
16.50 = ( 1 + g ) / ( 10% - g )
16.50 * ( 10% - g ) = 1 + g
1.65 - 16.50g = 1 + g
1.65 - 1 = g + 16.50g
0.65 = 17.50g
g = 0.65 / 17.50 3.71%